OpenStudy (anonymous):

1 OVER 5+√16 What does this simplify to?

OpenStudy (anonymous):

but wouldnt you multiply both sides not just by √16 but by 5+√16?

OpenStudy (campbell_st):

you need to evaluate the denominator \[\frac{1}{5 + \sqrt{16}} = \]

OpenStudy (campbell_st):

that results in rationalising the denominator and you actually multiply by \[\frac{5 - \sqrt{16}}{5 - \sqrt{16}}\] this will result in the denominator becoming the difference of 2 squares and hence a rational number \[\frac{1}{5 + \sqrt{16}} \times \frac{ 5 - \sqrt{16}}{5 - \sqrt{16}}\] which yeilds the same answer as shown above

OpenStudy (anonymous):

I am soo confused. how does 5-√16 Over 9 EQUAL 1/9?

OpenStudy (campbell_st):

ok so the denominator becomes 25 - 16 = 9 the numberator is \[1 \times ( 5 - \sqrt{16}) = 1 \times ( 5 - 4) = 1\] so you end up with 1/9

OpenStudy (campbell_st):

the denominator when multiplied is \[5^2 - (\sqrt{16})^2 = 25 - 16\]

OpenStudy (anonymous):

okay thanks, what about 3+√7 OVER 2-√10? Would it equal 5+√7 OVER -8?